OpenStudy (anonymous):

Solve the equation (4x + 1)(3x - 2) = 0 {-4, 1.5} {-1/4, 2/3} {1/4, 2/3} {-4, 3}

OpenStudy (anonymous):

if two numbers multiply to give zero, it means that one of the numbers must equal zero. a x b = 0, either a=0 or b=0 (or both could be zero). therefore in this case either 4x+1=0 or 3x-2=0. so solve for x in both cases to find the values of x.

OpenStudy (anonymous):

i don't get it D: